Two identical small metal coated spheres R and S are given identical charges of 40nC each. They are suspended at the same point from a ceiling by means of identical light inextesible insulating threads of equal length.the threads are of negligible mass. When the system is in equilibrium the angle between the threads is 60° while the distance between R and S is 5 cm. Calculate the mass of sphere R
1
Expert's answer
2020-12-14T07:16:33-0500


Find the electrostatic force F:

"F=k\frac{q^2}{r^2}."

The spheres are in equilibrium, so, the sum of forces along horizontal and vertical are zero. Begin with horizontal:


"F-T\text{ sin}30°=0,\\
T=\frac{F}{\text{sin}30°}."

Vertical:


"T\text{ cos}30°-mg=0, \\
m=\frac{T\text{ cos}30°}{g}=\frac{F}{g}\text{ cot}30°=\frac{kq^2}{gr^2}\text{ cot}30°=113\text{ mg}."